OpenStudy (anonymous):

What is the height of a cylinder with a volume of 1,899.7 cubic inches and a radius of 11 inches? (Use 3.14 for π.) 5 inches 6 inches 7 inches 8 inches

OpenStudy (reemii):

you need the formula for the volume of a cylinder. When you write it down, you will see that the statement gives the value of all quantities except "h" the height.

OpenStudy (anonymous):

what is the formula?

OpenStudy (reemii):

\(V = \pi r^2 h \) where \(V=\) volume, \(r=\) radius, \(h=\) height.

OpenStudy (reemii):

u can use a calculator? then 1889.7/(3.14 * 11^2) = 4.99.. = 5 (approx.)

OpenStudy (anonymous):

so would it be 502.4 ft^2 at least that is what i got

OpenStudy (reemii):

your answer is not for "h"..?

OpenStudy (anonymous):

sorry i was looking at the wrong problem sorry hold on

OpenStudy (reemii):

ok haha no problem.

OpenStudy (anonymous):

um 7 inches....? sorry i am really bad at math

OpenStudy (reemii):

Forom the formula above, if you want to find \(h\), use \(\frac{V}{\pi r^2} = h\). V = 1899.7, pi = 3.14, r = 11. the answer's not 7.

OpenStudy (reemii):

5 right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thank u so much for ur help!!!!

OpenStudy (reemii):

yw